# Can LLM “Self-report”?: Evaluating the Validity of Self-report Scales in Measuring Personality Design in LLM-based Chatbots

### v1

@@ Abstract
Personality design plays an important role in chatbot development. From rule-based chatbots to LLM-based chatbots, evaluating the effectiveness of personality design has become more challenging due to the increasingly open-ended interactions. A recent popular approach uses self-report questionnaires to assess LLM-based chatbots’ personality traits. However, such an approach has raised serious validity concerns: chatbot’s “self-report” personality may not align with human perception based on their interaction. Can LLM-based chatbots “self-report” their personality? We created 500 chatbots with distinct personality designs and evaluated the validity of self-reported personality scales in LLM-based chatbot’s personality evaluation. Our findings indicate that the chatbot’s answers on human personality scales exhibit weak correlations with both user perception and interaction quality, which raises both criterion and predictive validity concerns of such a method. Further analysis revealed the role of task context and interaction in the chatbot’s personality design assessment. We discuss the design implications for building contextualized and interactive evaluation of the chatbot’s personality design.
Can LLM “Self-report”?: Evaluating the Validity of Self-report Scales in Measuring Personality Design in LLM-based Chatbots
Huiqi Zou1, Pengda Wang2, Zihan Yan3, Tianjun Sun2, and Ziang Xiao1 1Department of Computer Science, Johns Hopkins University 2Department of Psychological Sciences, Rice University 3Cognitive Science and Language Processing Informatics, University of Illinois Urbana-Champaign {hzou11,ziang.xiao}@jhu.edu; {pw32,ts110}@rice.edu; zihan25@illinois.edu
@@ 1 Introduction
Figure 1: Overview of the evaluation pipeline. Two measures are used to evaluate the chatbot’s personality: (1) LLM self-reporting through established personality scales (e.g., Big Five Inventory-2 and IPIP-NEO-120); (2) Human-perceived personality, where users interact with the chatbot and subsequently rate its personality based on their interaction.
@@
With advancements in artificial intelligence (AI), especially in natural language processing (NLP), AI-based chatbots are now being used in various fields. Their applications, including creative writing (Chung et al., 2022), mental counseling (Lai et al., 2023), data collection (Wei et al., 2024), and teaching assistant (Hicke et al., 2023), have made interactions with conversational agents increasingly common. To further enhance dialogue engagement and adapt these agents to specific tasks, developers often manipulate chatbots’ behavior patterns or assign them specific identities. One common method is through personality design (Shao et al., 2023; Li et al., 2023; Park et al., 2023).
@@
The development of chatbot’s personality design has progressed significantly, from early rule-based systems such as ELIZA (Weizenbaum, 1966), which used predefined scripts to simulate human conversation, to modern methods involving large language models (LLMs). Initially, personality design was limited by rigid sets of rules and responses. However, as AI systems evolved, more flexible approaches emerged, such as end-to-end systems Zhang et al. (2018), which train the dialogue system to respond in a manner that aligns with a given persona. More recent developments center around using LLMs, which serve as the backbone of AI chatbots with customizable personalities. This includes characterizing LLM agents through demographic and persona-based dialogue, as well as using more dimensional approaches, such as the 70 bipolar adjectives proposed by Goldberg (1992); Serapio-García et al. (2023), or converting personality scale items into open-ended questions (Ran et al., 2024), offering greater flexibility in adapting to various chatbot personality designs. In this study, we will focus on LLM-based chatbots and their personality design.
As the chatbot becomes more open-ended, evaluating the effectiveness of the personality design becomes increasingly challenging. Unlike earlier systems, there are no hand-written rules for chatbot designers to inspect. However, the open-endedness of those chatbots also enables new evaluation methods. Researchers have started to use self-report personality scales designed for humans to evaluate the effectiveness of LLM-based chatbots’ personality design (Huang et al., 2023a; Wang et al., 2024c; Tu et al., 2023). In these studies, LLMs are typically asked to rate items on a Likert scale Likert (1932), with each item corresponding to a specific personality trait dimension. Subsequently, the overall personality scores in each domain are calculated based on these item-level ratings.
@@
While the “self-report” method can produce fast and low-cost evaluation results, there are strong, yet untested, assumptions regarding the validity of such methods. First, it assumes the internal consistency of LLMs, where the chatbot’s response on those scales is consistent with its behaviors, e.g., their interaction with people, which is the ultimate goal of evaluation. Second, these methods presume that the human experiences captured by psychometric scales also apply to LLMs. For instance, a chatbot designed for project management tasks might be assessed on extraversion using an item like “Love surprise parties” (Johnson, 2014a), which is irrelevant to the chatbot’s actual role in the designed task. Although these psychometric scales are well-validated for human samples, the examined validity does not directly transfer to the chatbot evaluation, raising concerns about the applicability of such methods in non-human settings.
@@
We aim to fill the gap and evaluate the validity of an LLM-based chatbot’s self-reported personality. Guided by measurement theory Xiao et al. (2023), we evaluated two sets of validity: Convergent and Discriminant validity, where we examined the intrinsic structure, and Criterion and Predictive validity, where we examined the extrinsic relationship with external variables. In particular, we looked at two external variables that are key to the ultimate goal of personality design: human perception and interaction quality.
@@
In this study, we created 500 chatbots with distinct personality designs and collected their “self-report” personality. Then, we recruited 500 participants to interact with each of the chatbots to complete the designed task, give their assessment of the chatbot’s personality, and rate the interaction quality. Figure 1 illustrates our evaluation pipeline. Our findings indicate that although the results of the “self-report” personality scales achieve moderate Convergent and Discriminant validity, they fail to align with human perception and exhibit weak correlations with interaction quality, which suggests limited Criterion and Predictive validity.
Our study raises significant validity concerns regarding the use of self-report personality scales for evaluating chatbot personality design. We argue that to design effective personality design evaluation methods, we need to consider how a chatbot’s personality manifests in its interaction with humans and ground such evaluation in downstream tasks. Future work should consider the rich signals encapsulated in human interaction.
Our work provides three contributions,
@@
  • Through an empirical study with 500 participants, we unveil the validity concerns of using self-report personality scales for evaluating LLM-based chatbot’s personality design.
@@
  • We present a dataset containing a rich log of human interactions with 500 chatbots, each with distinct personality designs, along with human perceptions of their personalities, facilitating the development of novel interaction-based personality evaluation methods.
@@ 2.1 Chatbot Design
In this study, we first created a set of LLM-based chatbots with various personality designs. Those chatbots were created for tasks in which personality design is critical in user experience. Our chatbot framework consists of two main components: a personality module (§2.1.1) and a task module (§2.1.2). The personality module shapes the chatbot’s behavior with predefined personality traits, while the task module provides role-specific instructions and task outlines to complete the task. The complete prompts are in Appendix A.
@@ 2.1.1 Personality Module
The personality design of our chatbot is based on the Big Five model John et al. (1999), a foundational framework in personality research that describes personality across five key domains: extraversion (EXT), agreeableness (AGR), conscientiousness (CON), neuroticism (NEU), and openness to experience (OPE). We adopted the shape approach from Serapio-García et al. (2023) to shape a chatbot’s personality in each domain. This method extends Goldberg (1992) list of 70 bipolar adjectives to 104 personality trait descriptors, which are mapped to different personality facets within each domain. For example, adjectives such as ‘unenergetic’ and ‘energetic’ represent the lower and higher levels of extraversion, respectively. These adjectives were then paired with linguistic qualifiers from Likert-type scales Likert (1932) to create personality profiles at varying levels. For our study, we selected the highest-level linguistic qualifier (i.e., ‘extremely’) and randomly sampled five high-level or low-level adjectives from the same personality domain to generate prompts.
@@ 2.1.2 Task Module
We selected five common chatbot task categories where personality design plays a critical role in user experience and task effectiveness. This approach allows us to evaluate the validity of the chatbot’s personality design via emulating real-world scenarios where chatbot designers rely on personality evaluation results to build an effective chatbot. The selected tasks are as follows:
@@
    Job Interview, where the chatbot works as an HR representative to assess the interviewee’s Organizational Citizenship Behaviors (OCB) such as initiative, helping, and compliance, was chosen due to the strong association of the interviewer’s conscientiousness in structured interview settings Heimann et al. (2021).
@@
    Public Service, focusing on handling workday challenges like resource allocation and crisis management, highlights the chatbot’s ability to help humans navigate high-stress situations where agreeableness and emotional stability are crucial Mishra et al. (2023).
@@
    Personalized Social Support, where chatbot features are rooted in the counseling psychology principles of empathy and reflective listening as outlined by Rogers (1957), offers empathetic responses to individuals in distress referring Acceptance and Commitment Therapy (ACT) techniques. Research shows that empathy is closely linked to agreeableness Walton et al. (2023).
@@
    Customized Travel Planning, which relies on the chatbot’s decision-making and personalization skills, underscores the importance of anthropomorphic traits in recommendation systems that foster trust and enhance user decision-making Qiu and Benbasat (2009). Additionally, previous research suggests that question types related to extraversion levels significantly impact user satisfaction Miyama and Okada (2022).
@@
    Guided Learning, where the chatbot aids humans in understanding complex topics, is based on the constructivist learning approach Higgs et al. (2004) and highlights the critical role of extraversion in instructors for enhancing student acceptance.
In total, we randomly sampled five personality adjective markers at high and low levels of each domain ten times. Combining these 100 personality descriptions with five task instructions generated 500 chatbots with distinct personality designs. We used GPT-4o OpenAI (2024), one of the most performant models available, as the backbone model with the temperature set to zero.
@@ 2.2 Personality Design Evaluation
2.2.1 Self-reported Personality
To quantify the chatbot’s self-reported personality, we utilized three well-established psychometric tests that assess the Big Five domains John et al. (1999): Big Five Inventory–2 Extra-Short Form (BFI-2-XS) Soto and John (2017b), BFI-2 Soto and John (2017a) and IPIP-NEO-120 Johnson (2014b). The chatbot’s responses to test items were aggregated to calculate the personality scores for the five domains respectively, which were then used for subsequent analysis. Following the work of Serapio-García et al. (2023), we instructed the chatbot to rate test items using a standardized response scale from one to five based on the personality descriptions. The complete prompt format is shown in Appendix B.
@@ 2.2.2 Human-perceived Personality
We conducted a human study to gather participants’ perceptions of the chatbot’s personality. Each participant was randomly assigned to interact with one chatbot and engaged in multi-turn, text-based conversations to complete the designed task. A user interface example is shown in Appendix C.
After completing the task, participants were asked to rate their perceived personality of the chatbot using the BFI-2-XS Soto and John (2017b). The BFI-2-XS consists of 15 items, each representing a distinct facet of one of the Big Five personality domains, thus preserving the scale’s comprehensive descriptive and predictive capabilities.
@@ 2.3 Study Procedure
To gather the human-perceived personality of each chatbot, we recruited English-speaking participants from Prolific [^1] https://www.prolific.com, with each participant assessing one chatbot. At the beginning of the survey, participants were given the task’s objective, along with clear instructions on how to start the conversation, including an example to clarify how the participant might approach the task. A privacy reminder is included to ensure no private information is shared. After interacting with the chatbot, each participant filled out the chatbot personality evaluation questionnaire and completed a separate User Experience Questionnaire (UEQ) Laugwitz et al. (2008). The entire process was designed to take 10 to 15 minutes. Appendix F presents detailed participant statistics and their demographic information.
2.4 Construct Validity Evaluation
Construct validity refers to the extent to which a test accurately measures what it is intended to measure Cronbach and Meehl (1955). In our study, we assessed validity through four aspects:

### v2

@@ Abstract
A chatbot’s personality design is key to interaction quality. As chatbots evolved from rule-based systems to those powered by large language models (LLMs), evaluating the effectiveness of their personality design has become increasingly complex, particularly due to the open-ended nature of interactions. A recent and widely adopted method for assessing the personality design of LLM-based chatbots is the use of self-report questionnaires. These questionnaires, often borrowed from established human personality inventories, ask the chatbot to rate itself on various personality traits. Can LLM-based chatbots meaningfully “self-report” their personality? We created 500 chatbots with distinct personality designs and evaluated the validity of their self-report personality scores by examining human perceptions formed during interactions with these chatbots. Our findings indicate that the chatbot’s answers on human personality scales exhibit weak correlations with both human-perceived personality traits and the overall interaction quality. These findings raise concerns about both the criterion validity and the predictive validity of self-report methods in this context. Further analysis revealed the role of task context and interaction in the chatbot’s personality design assessment. We further discuss design implications for creating more contextualized and interactive evaluation.
@@ 1 Introduction
With recent advancements in artificial intelligence (AI), particularly in natural language processing (NLP), AI-based chatbots have found applications across various fields. Their use in areas, such as creative writing (Chung et al., 2022), mental health counseling (Lai et al., 2023), data collection (Wei et al., 2024) and educational support as teaching assistants (Hicke et al., 2023), has made interactions with conversational agents increasingly common. To further enhance dialogue engagement and tailor these agents to specific tasks, developers often manipulate chatbots’ behaviors or assign them particular identities. One prevalent strategy is through personality design (Shao et al., 2023; Li et al., 2023; Park et al., 2023).
Figure 1: Overview of the Evaluation Pipeline. Two measures are used to evaluate the chatbot’s personality: (1) Chabot self-reporting through established personality scales (e.g., BFI-2 and IPIP-NEO-120); (2) Human-perceived personality, where human interact with the chatbot and subsequently rate its personality based on their interaction.
@@
The design of chatbot personalities has evolved significantly, evolving from early rule-based systems such as ELIZA (Weizenbaum, 1966), which relied on predefined scripts to simulate human conversation, to contemporary approaches based on large language models (LLMs). The involvement of LLMs enables various methods of personality instantiation, such as characterizing agents through demographic and persona-based dialogue, as well as employing more fine-grained approaches. For instance, Goldberg (1992)’s set of 70 bipolar adjectives, mapped to the Big Five personality domains (Serapio-García et al., 2023), and their facets have been used to characterize chatbot behaviors. Similarly, researchers have transformed personality scale items into open-ended prompts to elicit personality-consistent responses from LLMs (Ran et al., 2024).
As chatbots become increasingly open-ended in their capabilities, evaluating the effectiveness of their personality design presents growing challenges. Unlike earlier rule-based systems, modern chatbot architectures lack hand-crafted rules that designers can easily examine or adjust. However, this openness also facilitates the development of novel evaluation methods. Recent research has begun to explore the use of self-report personality scales, originally developed for assessing human traits, to evaluate the personality constructs of LLM-based chatbots (Huang et al., 2023a; Wang et al., 2024c; Tu et al., 2023). In these studies, LLMs are typically prompted to respond to items from standardized personality inventories using a Likert scale format (Likert, 1932), where each item works as an indicator to reflect a specific personality construct dimension. Responses to individual items are summarized to compute aggregate scores for each personality domain.
@@
Although the “self-report” method is convenient and cost-effective for evaluating LLMs, it relies on assumptions that have not yet been systematically validated through empirical research. First, it assumes a high degree of alignment between the personality traits attributed to an LLM-based chatbot by design and those perceived by humans during interactions (the ultimate goal of evaluation). However, even in human psychological research, the consistency between self-reports and informant reports tends to be limited due to differences in informational sources (e.g., Connolly et al., 2007; Connelly & Ones, 2010; McCrae & Costa, 1987). Second, these methods presume that the structure and content of psychological assessment tools developed for humans can also be applied to LLM-based chatbots. For instance, a chatbot designed for project management might be assessed on Extraversion using an item like “Love surprise parties” (Johnson, 2014a), which is irrelevant to the chatbot’s actual role in the designed task. Although these psychometric scales are well-validated for human samples, the examined validity does not directly transfer to chatbot evaluation, raising concerns about the applicability of such methods in non-human settings.
@@
This study aims to fill the gap by evaluating the validity of an LLM-based chatbot’s self-report personality. Guided by principles from classical test theory and measurement theory (e.g., Allen & Yen, 2001; Xiao et al., 2023), we evaluated two sets of validity: Convergent and Discriminant validity, where we examined the intrinsic structure, and Criterion and Predictive validity, where we examined the extrinsic relationship with relevant variables. In particular, we focus on two external variables central to the ultimate goal of personality design in AI systems: human perception and interaction quality.
@@
In this study, we created 500 chatbots, each with distinct personality designs, and collected their “self-report” personality. We recruited 500 human participants, each assigned to interact with one of the chatbots to complete a designated task, evaluate the chatbot’s personality, and rate the interaction quality. Figure 1 presents an overview of the evaluation pipeline. Our findings indicate that while the self-report personality scales demonstrate moderate Convergent and Discriminant validity, they show poor alignment with human-perceived personality scores and weak correlations with the human-rated interaction quality, suggesting limited Criterion and Predictive validity.
Our study reveals the limitations of relying solely on self-report measures to assess personality design in LLM-based chatbots. We argue that effective evaluation methods must account for how a chatbot’s personality emerges through interaction with human users and should be grounded in task-based performance. Future research should consider the rich signals in human-chatbot interactions and incorporate human perception as a critical dimension in evaluating personalized AI systems. Our work makes three main contributions:
@@
  • Through an empirical study with 500 participants, we highlight the validity concerns of using self-report personality scales to evaluate the personality design of LLM-based chatbots.
@@
  • We present a dataset containing a rich log of human interactions with 500 chatbots, each with distinct personality designs and human perceptions of their personalities, facilitating the development of interaction-based personality evaluation methods.
2 Related Work
Human-Like Responses in LLMs.
Recent studies have shown that LLMs demonstrate remarkable abilities resembling human-like responses. For example, Lampinen et al. (2024) found that LLMs display response patterns similar to humans in logical reasoning tasks, with similar findings being reported by Kosinski (2023) and Wang et al. (2024a). Additionally, Pellert et al. (2023) found that LLMs perform comparably to humans across various psychological scales, such as value orientations and moral norms. All these abilities have garnered significant interest from both computer scientists and social scientists, as these findings suggest that we may be able to draw upon existing social science research on humans as a reference for LLM studies. For instance, personality, which has been extensively studied in psychology and consistently proven to be a reliable predictor of various behavioral and psychological outcomes, is one such area of interest. Personality encodes rich and complex information in textual data  (Goldberg, 1990; Saucier & Goldberg, 2001). LLMs may capture and model such encoding by learning from vast training data. Many studies explore the personality traits LLMs exhibit, with the hope to leverage these simulated traits for predicting or shaping model behavior (Lee et al., 2024; Huang et al., 2023b; Wang et al., 2024b; 2025; Ran et al., 2024).
Current Approaches for Measuring LLM-based Chatbots’ Personality.
Currently, most assessments of LLM-based chatbot personality rely on psychological frameworks developed for human personality measurement, typically employing standardized personality scales. A common approach is to have the model respond to scales that have already been developed and validated on human samples, such as the Big Five Inventory-2 (BFI-2) (Soto & John, 2017a) and HEXACO (Lee & Ashton, 2004; 2006). The model’s scores are then calculated based on its answers to the multiple-choice questions (MCQs) on these scales. The rationale behind this is straightforward: since we believe that LLMs simulate human abilities, we can directly borrow well-established and validated methods used for humans to assess the personality traits of these models. For example, studies like Huang et al. (2023b) used this approach to evaluate whether LLMs reflect the personality traits outlined in the BFI (John et al., 1999). Serapio-García et al. (2023) presented the reliability and validity of LLMs in resembling personality profiles along desired dimensions with International Personality Item Pool– Neuroticism, Extraversion, Openness (IPIP-NEO) (Goldberg et al., 1999) and BFI (John et al., 1999). However, some researchers argue that traditional scales may not sufficiently capture the personality of LLMs and have adopted alternative approaches. For instance, some researchers modified scale questions into prompts to elicit text-based responses (Wang et al., 2024c), while others have experts rate stories generated by LLMs according to specified criteria (Jiang et al., 2023). These evaluation methods may more accurately reflect real-world application scenarios.
Self-Reports vs. Informant Reports.
Self-report methods have limitations in capturing the behavior nuances that reflect personality traits of LLM-based chatbots, as they are not grounded in user experience. Given that the primary goal of chatbot personality design is to enhance engagement and provide personalized interactions, assessment methods should center on human perception. This challenge parallels long-standing exploration in psychology regarding the distinction between self-reports and informant reports (e.g., Connolly et al., 2007; Connelly & Ones, 2010; McCrae & Costa, 1987). While self-reports capture an individual’s subjective understanding of their own traits, informant reports offer an external perspective based on others’ interpretations of the individual’s behavior. A meta-analysis in psychology found that the average correlation between self-reports and informant reports for the Big Five personality traits was only 0.36 (Connolly et al., 2007), indicating overlap and divergence. For chatbots, task-specific design further influences personality expression, potentially widening the gap between their self-report and human-perceived traits. Thus, relying solely on self-reports may fail to reflect how chatbots actually affect user experience.
@@ 3.1 Chatbot Design
In this study, we first created a set of LLM-based chatbots with various personality designs. Those chatbots were created for tasks in which personality design is critical for user experience. Our chatbot framework consists of two main components: a personality module (§3.1.1) and a task module (§3.1.2). The personality module shapes the chatbot’s behavior with predefined personality traits, while the task module provides role-specific instructions and task outlines to complete the task. The complete prompts are in Appendix A.
@@ 3.1.1 Personality Module
The personality design of our chatbot is based on the Big Five model (John et al., 1999), a foundational framework in personality research that describes personality across five key domains: extraversion (EXT), agreeableness (AGR), conscientiousness (CON), neuroticism (NEU), and openness to experience (OPE). We adopted the shape approach from Serapio-García et al. (2023) to shape a chatbot’s personality in each domain. This method extends Goldberg (1992) list of 70 bipolar adjectives to 104 personality trait descriptors, which are mapped to different facets within each domain. For example, adjectives such as ‘unenergetic’ and ‘energetic’ represent the lower and higher levels of Extraversion, respectively. These adjectives were then paired with linguistic qualifiers from Likert-type scales (Likert, 1932) to create personality profiles at varying levels. For our study, we selected the strongest linguistic qualifier (i.e., ‘extremely’) and randomly sampled five high-level or low-level adjectives from the same personality domain to generate prompts. Each chatbot is prompted to exhibit either a high or low level of a single personality trait, with others left unconstrained.
@@ 3.1.2 Task Module
We selected five chatbot task categories where personality design plays a critical role in user experience and task effectiveness. This approach allows us to evaluate the validity of the chatbot’s personality design via emulating real-world scenarios where designers rely on personality evaluation results to build an effective chatbot. The selected tasks are as follows:
@@
    Job Interview, where the chatbot works as an HR representative to assess the interviewee’s Organizational Citizenship Behaviors (OCB), such as initiative, helping, and compliance, was chosen due to the strong association of the interviewer’s conscientiousness in structured interview settings (Heimann et al., 2021).
@@
    Public Service, focusing on handling workday challenges like crisis management, highlights the chatbot’s ability to help humans navigate high-stress situations where agreeableness and emotional stability are crucial (Mishra et al., 2023).
@@
    Personalized Social Support, where chatbot features are rooted in the counseling psychology principles of empathy and reflective listening as outlined by Rogers (1957), offers empathetic responses to individuals in distress referring Acceptance and Commitment Therapy (ACT) techniques. Research shows that empathy is closely linked to agreeableness (Walton et al., 2023).
@@
    Customized Travel Planning, which relies on the chatbot’s decision-making and personalization skills, underscores the importance of anthropomorphic traits in recommendation systems that foster trust and enhance user decision-making (Qiu & Benbasat, 2009). Previous research also suggests that question types related to extraversion levels significantly impact user satisfaction (Miyama & Okada, 2022).
@@
    Guided Learning, where the chatbot aids humans in understanding complex topics, is based on the constructivist learning approach (Higgs et al., 2004) and highlights the critical role of extraversion in instructors for enhancing student acceptance.
In total, we randomly sampled five personality adjective markers at high or low levels of each domain ten times, respectively. Combined with five task instructions, this resulted in 500 distinct chatbot personality designs. We used GPT-4o (OpenAI, 2024), one of the most performant models available, as the backbone model with the temperature set to zero.
@@ 3.2 Personality Design Evaluation
3.2.1 Self-report Personality
To quantify the chatbot’s self-report personality, we utilized three well-established psychometric tests that assess the Big Five domains (John et al., 1999): Big Five Inventory–2 Extra-Short Form (BFI-2-XS) (Soto & John, 2017b), BFI-2 (Soto & John, 2017a) and IPIP-NEO-120 (Johnson, 2014b). The chatbot’s responses to test items were aggregated to calculate the personality scores for the five domains respectively, which were then used for subsequent analysis. Following the work of Serapio-García et al. (2023), we instructed the chatbot to rate test items using a standardized response scale from one to five based on the personality descriptions. The complete prompt format is shown in Appendix B.
@@ 3.2.2 Human-perceived Personality
We conducted a human study to gather participants’ perceptions of the chatbot’s personality with the designed interface (Appendix C). Each participant was randomly assigned to interact with one chatbot and engaged in multi-turn, task-based conversations.
After completing the task, participants were asked to rate their perceived personality of the chatbot using the BFI-2-XS (Soto & John, 2017b). The BFI-2-XS consists of 15 items, each representing a distinct facet of one of the Big Five personality domains, thus preserving the scale’s comprehensive descriptive and predictive capabilities. To mitigate the impact of individual variability, we aggregated evaluation scores across chatbots sharing the same personality configuration (i.e., for all high or all low) for analysis.
@@ 3.3 Study Procedure
To gather the human-perceived personality of each chatbot, we recruited English-speaking participants from Prolific [^1] https://www.prolific.com, with each participant assessing one chatbot. At the beginning of the survey, participants were given the task’s objective, along with clear instructions on how to start the conversation, including an example to clarify how the participant might approach the task. A privacy reminder is included to ensure no private information is shared. After interacting with the chatbot, each participant filled out the chatbot personality evaluation questionnaire and completed a separate User Experience Questionnaire (UEQ) (Laugwitz et al., 2008). The entire process was designed to take 10 to 15 minutes. Appendix E presents a summary of the collected dataset statistics. Appendix F presents detailed participant statistics and their demographic information.
3.4 Analysis Method
Our analysis evaluates the effectiveness of personality settings and the alignment between self-reports and human perceptions using descriptive statistics and correlation analyses. We began by computing the means and standard deviations of both self-report and human-perceived personality scores to examine overall distribution patterns. Statistical significance tests were conducted to evaluate differences between the high and low personality conditions, as well as between human perceptions and the chatbot’s self-report scores. In addition, we assessed construct validity, the extent to which a test accurately measures what it is intended to measure (Cronbach & Meehl, 1955), based on four psychometric dimensions: